How much heat is required to convert 20.0 [g] of ice at −10.0 ∘C to steam at 100.0 ∘C?


1
Expert's answer
2021-02-03T16:12:00-0500
"Q=Q_1+Q_2+Q_3+Q_4,""Q=m_ic_i\Delta T+m_iL_f+m_wc_w\Delta T+m_wL_v."

Let's find the amount of heat required to convert 0.02 kg of ice at -10°C to 0.02 kg of ice at 0°C:


"Q_1=m_ic_i\Delta T=0.02\ kg\cdot2100\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot10^{\circ}C=420\ J."

Let's find the amount of heat required to convert 0.02 kg of ice at 0°C to 0.02 kg of water at 0°C:


"Q_2=m_iL_f=0.02\ kg\cdot3.34\cdot10^5\ \dfrac{J}{kg}=6680\ J."

Let's find the amount of heat required to convert 0.02 kg of water at 0°C to 0.02 kg of water at 100°C:


"Q_3=m_wc_w\Delta T=0.02\ kg\cdot4180\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot100^{\circ}C=8360\ J."


Let's find the amount of heat required to convert 0.02 kg of water at 100°C to 0.02 kg of steam at 100°C:


"Q_4=m_wL_v=0.02\ kg\cdot2.264\cdot10^6\ \dfrac{J}{kg}=45280\ J."

Finally, the total amount of the heat required to convert 0.02 kg of ice at -10°C to steam at 100°C:


"Q=420\ J+6680\ J+8360\ J+45280\ J=60740\ J."